### v1

@@ Abstract
We present latentSplat, a method to predict semantic Gaussians in a 3D latent space that can be splatted and decoded by a light-weight generative 2D architecture. Existing methods for generalizable 3D reconstruction either do not enable fast inference of high resolution novel views due to slow volume rendering, or are limited to interpolation of close input views, even in simpler settings with a single central object, where 360-degree generalization is possible. In this work, we combine a regression-based approach with a generative model, moving towards both of these capabilities within the same method, trained purely on readily available real video data. The core of our method are variational 3D Gaussians, a representation that efficiently encodes varying uncertainty within a latent space consisting of 3D feature Gaussians. From these Gaussians, specific instances can be sampled and rendered via efficient Gaussian splatting and a fast, generative decoder network. We show that latentSplat outperforms previous works in reconstruction quality and generalization, while being fast and scalable to high-resolution data. Our project website111geometric-rl.mpi-inf.mpg.de/latentsplat/ provides additional results including videos.
@@ 1 Introduction
Performing 3D reconstruction from a single or a few images is a longstanding goal in computer vision, which went through many iterations of advancements, most recently driven by new techniques in the areas of generalizable radiance fields [53, 10] and foundational diffusion models [40, 35]. At the core, the task is to find an optimal 3D representation that fits a set of observations, a task that is highly underconstrained as there usually are an infinite amount of valid reconstructions that satisfy the given observations. Thus, a strong prior is needed to find a fitting solution - usually modeled by deep neural networks trained on a large amount of data. When designing methods to learn data priors for 3D reconstruction, efficiency is a crucial aspect to allow training on large datasets required for generalization. In this work, we present a method that is highly efficient, scales to a large amount of data and can be trained on real video data without 3D supervision. Real video data is readily available in vast quantities and a promising data type for large 3D models.
Recently, there have been two lines of solutions for the given task of generalizable reconstruction: regression-based approaches and generative approaches. Methods based on regression, such as pixelNeRF [53] or pixelSplat [9], are usually efficient but are only trained to predict the mean of all possible solutions. While they often succeed in predicting high quality reconstructions of regions that strongly correlate with the input observations, they struggle in regions of high uncertainty, collapsing to blurry reconstructions that lack high frequency details or fail entirely in unseen areas of larger scenes.
@@
Regression-based models have often been used as conditioning for generative models. A well-known, traditional example are Variational Autoencoders (VAEs) [23]. An encoder is used to parameterize a variational distribution in latent space, from which we can sample a specific latent vector that can be decoded into an element following the data distribution via a decoder. It is promising to bring this concept efficiently to 3 dimensions, where a regression model is used to estimate the uncertainty for different locations in 3D space individually, providing the desired locality in uncertainty modeling.
@@
  • outperforms recent methods in two view reconstruction, achieving state-of-the-art quality both quantitatively and qualitatively, especially in challenging cases of wide-spread input views and view extrapolation,
@@
  • is applicable to both, object-centric (with 360° NVS) and general scenes,
@@ 2 Related Work
We revisit recent methods in the area of generalizable novel view synthesis (NVS) and 3D reconstruction. Neural fields [33, 29, 5, 43] have been the most dominant representation to store 3D information for single scenes and scene/object data distributions. Lately, more explicit representations live through a renaissance with 3D Gaussian Splatting [22]. Due to the challenging nature of the task, most methods that enable 360 degree generalization only perform reconstruction on the level of individual objects. Even with the recent rise of large-scale generative models, methods that perform extrapolating reconstruction of larger scenes are still a rarity, often due to missing scalability of the methods. In the following, we distinguish between regression-based approaches in Sec. 2.1 and generative approaches in Sec. 2.2.
@@ 2.1 Regression-based Generalizable NVS
Several regression-based models that perform reconstruction from a few views have been proposed in the recent years. A large line of work performs generalization over object categories [45, 19, 15, 51] and are not capable of generalization on scenes. While some early methods in that direction are theoretically capable of generalization on a scene level they often fail to provide the necessary capacity or efficiency [53, 12, 24, 31]. For larger scenes, several image-based rendering methods have been proposed [1, 36, 39, 52]. They produce high quality results in view interpolation settings but cannot generalize to unseen areas. A related approach is to predict multi-plane images [46, 49, 55, 56]. Since generalization is limited to multiple planes in 3D these methods do only allow small view-point variations. Multi-view stereo is also a popular way to provide geometry priors for novel-view synthesis with deep learning [10, 39], which has been modeled fully with deep learning architectures as well [20]. Several alternative representations have been introduced to perform novel view synthesis, such as neural rays [27], light fields [44, 48], and patches [47].
In contrast to all of the above, we are able to provide high-quality 360° reconstructions of object-centric scenes as well as view inter- and extrapolation on large scenes, given only 2 input views. Closest to our work is the very recent pixelSplat [9]. In contrast to their purely regression-based approach, we (1) introduce a semantic feature representation instead of purely explicit Gaussians and (2) model uncertainty explicitly in our framework and therefore enable correct generalization to out-of-context views. Thus, we are able to reconstruct full objects in high quality, even if they are only partially observed with two views.
@@ 2.2 Generative Models for NVS
Generative approaches succeed in situations with high uncertainty, i.e. when the signal coming from the conditioning is not sufficient to determine the full reconstruction, or if there is no conditioning at all. A large line of work performs 3D generation and reconstruction of objects with 3D-aware GAN architectures by sampling or inversion, respectively [6, 34]. While these methods are able to produce high-quality results, they are not applicable to scenes. Similarly, 2D diffusion models have been recently adapted to perform the task of 3D reconstruction of objects [26, 25, 50, 57]. Another line of research trains diffusion models directly on 3D representations such as voxel grids [30, 21], triplanes [2, 11], or point clouds [42, 28], and can implement 3D reconstruction by guiding the diffusion process with gradients from reconstructing input views. However, these methods are limited to objects and do not scale to large scenes. Autoregressive transformers have been shown to be able to synthesize novel views that are consistent to some extent with the past sequence of views, but they fail to successfully leverage explicit 3D biases [41, 38].
One of the first works that applies 2D diffusion on larger scenes is GeNVS [8], which is also closest to our work. GeNVS [8] conditions an image space diffusion model with a pixelNeRF feature volume. In contrast to their approach, our approach is orders of magnitude faster and can scale easier to larger images, due to the efficient Gaussian representation instead of volume rendering, and a lightweight decoder instead of an expensive diffusion sampling procedure.
Figure 2: We present latentSplat, a method for scalable generalizable 3D reconstruction from two views. The architecture follows an autoencoder structure. (Left) Two input reference views are encoded into a 3D variational Gaussian representation using an epipolar transformer and a Gaussian sampling head. (Center) Variational Gaussians allow sampling of spherical harmonics feature coefficients that determine a specific instance of semantic Gaussians. (Right) The sampled instance can be rendered efficiently via Gaussian splatting and a light-weight VAE-GAN decoder.
@@ 3.3 Encoding Reference Views
The variational 3D Gaussian representation described in the previous section is obtained from two given reference views $\mathbf{I}_{1}$, $\mathbf{I}_{2}$ sampled from a video sequence. To this end, we adapt the epipolar transformer encoder from pixelSplat [9] to our setting of variational Gaussians by adding the capability of predicting means $\mathbf{h}_{\sigma}$ and variances $\mathbf{h}_{\sigma}$ of spherical harmonic coefficients for view-dependent features of each predicted 3D Gaussian (c.f. Fig 2, left). The encoder consists of three parts: (1) a vision transformer [32], (2) an epipolar transformer [16], and (3) a per-pixel sampling of 3D Gaussians from predicted depth distributions [9], which are shortly outlined in the following. For more details about the encoder architecture, we refer to the supplementals, Charatan et al. [9], and He et al. [16].
@@
With the described encoder, uncertainty in reconstruction is modeled in two ways: First, the Gaussian locations are sampled from the predicted distributions over the rays, modeling uncertainty in 3D Gaussian location. Second, uncertainty in local appearance is modeled via distribution parameter prediction of the variational Gaussians.
@@ 3.5 Training
Reconstruction Losses Similar to LDM [40], we use a combination of a standard L1 loss and LPIPS as a perceptual loss between the decoded feature image $\hat{\mathbf{T}}$ and the target image $\mathbf{T}$:
@@
Auxiliary Losses We further apply an auxiliary loss directly between the color renderings $\hat{\mathbf{T}}_{\mathrm{aux}}$ and the target image $\mathbf{T}$ to provide better gradients to the structural parameters (c.f. Sec 3.2):
@@
Generative Loss To enable correct sampling and generation in uncertain regions, we further optimize our method using GAN losses, adapted from the LDM VAE-GAN decoder [40]. We add and further train a pre-trained discriminator network [18] $D$ from LDM[40] predicting the likelihood of image patches being real. Therefore, our architecture is trained to directly maximize its output:
@@ 4 Experiments
In this section, we detail our experiments made with latentSplat, provide comparisons with state-of-the-art baselines, and verify our architecture design in form of an ablation. Our experiments aim to support the following statements: (1) latentSplat improves on previous methods for two-view interpolation in terms of visual quality, (2) can also generalize better to extrapolation, i.e. producing novel views outside of given reference views, (3) avoids unnecessary hallucination but sticks to the identity of the observed scene, and (4) maintains the real-time rendering capabilities and memory efficiency of 3D Gaussian splatting. We provide additional results in the appendix.
@@ 4.1 Experimental Setup
Datasets We conduct two-view reconstruction experiments for an object-centric setting as well as for general video datasets capturing diverse scenes. For the former, we make use of the Common Objects in 3D (CO3Dv2) [37] dataset, which consists of video captures of real-world objects grouped into categories. Following related work [8], we choose cleaned subsets of hydrants and teddybears, and randomly split the scenes into 95% training and 5% test data. In all experiments, including baselines, we gradually increase the gap between the two reference views from initially 8-18 frames up to 25 frames roughly corresponding to 90° in the usually 102 frames long sequences. At the same time, we increasingly randomize the target view selection from pure interpolation until a uniform distribution over all views. For an evaluation on general scenes, we test latentSplat on RealEstate10k [56], a dataset of home walkthrough clips gathered from about 10000 YouTube videos. We use the provided splits and employ the same training curriculum as pixelSplat [9] except of increasing the sampling interval of target views up to 45 frames before and after both reference views to enable learning extrapolation.
Baselines We compare our approach against four baselines. pixelNeRF [53] conditions a single NeRF MLP by interpolating pixel-aligned features of reference views. Du et al. [14] is a light field rendering approach that uses multi-view self-attention and cross-attention of target rays to samples along its epipolar lines in the input images. pixelSplat[9] defines an encoder including an epipolar transformer to predict 3D Gaussians along the rays of two reference views that can be efficiently rendered via rasterization. Unlike the previous regression-based approaches, GeNVS [7] proposes a generative diffusion model with view-conditioning via pixelNeRF in a feature space.
@@
Teddybears
Ours 53.47 0.004 0.338 0.173 20.83 0.663 71.12 0.010 0.434 0.219 17.71 0.533
Figure 3: Qualitative results on the CO3D dataset [37]. We evaluate on two categories. The top three rows depict the rendering of the target view on Hydrants, while the bottom three rows showcase the corresponding rendering of the target view on Teddybear, in comparison with the baselines.
Metrics We employ three groups of each two metrics: (1) FID [17] and KID [3] measure the similarity between the distributions of predicted novel views and the corresponding ground truth. They are the established metrics for image synthesis of generative models and reflect visual quality. (2) Perceptual metrics like LPIPS [54] and DISTS [13] leverage features of deep networks for the comparison of images w.r.t. structure and texture. (3) Classical reconstruction metrics such as PSNR and SSIM are still well-established for evaluation of dense-view 3D reconstruction. However, because incorporating plain pixel-wise similarities, these metrics prefer blur over realistic details and are therefore not well-suited for the evaluation of generative methods.
Further implementation details can be found in the supplementary materials and in our code, which we will make publicly available.
@@ 4.3 Scene-Level 3D Reconstruction
We report quantitative results for scene-level 3D reconstruction on RealEstate10k in Table 2. Again, latentSplat outperforms all baselines in FID and KID as well as LPIPS and DISTS. Although extrapolation on scene-level is a quite different task compared to learning a category-level prior for object-centric videos, we observe the same result that latentSplat generalizes better to extrapolation compared to all baselines, even achieving state-of-the-art in PSNR as well. This highlights the applicability of our method for various kinds of real-world videos. Looking at qualitative examples in Fig. 4, we can see that our approach produces clean and visually pleasing novel views with significantly less artifacts compared to the baselines.
@@
Figure 4: Qualitative results on RealEstate10k [56]. We show that in many challenging cases latentSplat provides better reconstruction than previous approaches. This specifically holds when (1) reference views are far apart from each other or when (2) extrapolating outside of reference views.
4.4 Comparison with GeNVS
We provide an additional qualitative comparison with GeNVS [8] in Fig. 5. Their code is not available and we are not able to reproduce their results for a fair comparison. We still aim to do a comparison using figures from their paper. Note that there are major differences between the setups of GeNVS and latentSplat such as the number of input views and the version of the dataset (CO3Dv1 vs CO3Dv2). We match their setting as close as possible by creating a test split of all their evaluation examples, finding closest poses, and using the same input view together with a close second view. Although the comparison of single- and two-view 3D reconstruction is not fair, note that GeNVS uses scale normalization and depth cues (c.f. supplementary material C.3) to resolve the scale ambiguity from a single input view. The qualitative comparison shows similar image quality. However, our approach is much faster and efficient during both inference and training, as detailed in Sec. 4.5.
Figure 5: Comparison with GeNVS [8] We qualitatively compare against GeNVS. Note that the setups of both methods differ strongly and that GeNVS code is not available for reproducing results. Thus, this comparison is not fair. We selected the same CO3D test examples and compared against renderings shown in their paper. Overall, both methods generate similar quality while ours is much faster (c.f. Sec. 4.5).

### v2

@@ Abstract
We present latentSplat, a method to predict semantic Gaussians in a 3D latent space that can be splatted and decoded by a light-weight generative 2D architecture. Existing methods for generalizable 3D reconstruction either do not scale to large scenes and resolutions, or are limited to interpolation of close input views. latentSplat combines the strengths of regression-based and generative approaches while being trained purely on readily available real video data. The core of our method are variational 3D Gaussians, a representation that efficiently encodes varying uncertainty within a latent space consisting of 3D feature Gaussians. From these Gaussians, specific instances can be sampled and rendered via efficient splatting and a fast, generative decoder. We show that latentSplat outperforms previous works in reconstruction quality and generalization, while being fast and scalable to high-resolution data. Our project website111geometric-rl.mpi-inf.mpg.de/latentsplat/ provides additional results, videos, and code.
@@ 1 Introduction
Performing 3D reconstruction from a single or a few images is a longstanding goal in computer vision, which went through many iterations of advancements, most recently driven by new techniques in the areas of generalizable radiance fields [76, 11] and foundational diffusion models [40, 46, 52, 50, 44]. At the core, the task is to find an optimal 3D representation that fits a set of observations, a task that is highly underconstrained as there are usually an infinite amount of valid reconstructions that satisfy the given observations. Thus, a strong prior is needed to find a fitting solution - usually modeled by deep neural networks trained on a large amount of data. When designing methods to learn data priors for 3D reconstruction, efficiency is a crucial aspect to allow training on large datasets required for generalization. In this work, we present a method that is highly efficient, scales to a large amount of data and can be trained on real video data without 3D supervision. Real video data is readily available in vast quantities and a promising data type for large 3D models.
Recently, there have been two lines of solutions for the given task of generalizable reconstruction: regression-based approaches and generative approaches. Methods based on regression, such as pixelNeRF [76] or pixelSplat [10], are usually efficient but are only trained to predict the mean of all possible solutions. While they often succeed in predicting high-quality reconstructions of regions that strongly correlate with the input observations, they struggle in regions of high uncertainty, collapsing to blurry reconstructions that lack high frequency details or fail entirely in unseen areas of larger scenes.
@@
Regression-based models have often been used as conditioning for generative models. A well-known example are Variational Autoencoders (VAEs) [32]. An encoder is used to parameterize a variational distribution in latent space, from which we can sample vectors that can be decoded into elements following the data distribution. It is promising to bring this concept efficiently to 3 dimensions, where a regression model estimates the uncertainty for different locations in 3D space individually, providing the desired locality in uncertainty modeling.
@@
  • outperforms recent methods in two-view reconstruction, achieving state-of-the-art quality both quantitatively and qualitatively, especially in challenging cases of wide-spread input views and view extrapolation,
@@
  • is applicable to both, object-centric (with 360° views) and general scenes,
  • enables downstream mesh reconstruction via 3D consistent novel views,
@@ 2 Related Work
We revisit recent methods in the area of generalizable novel view synthesis (NVS) and 3D reconstruction. Neural fields [37, 42, 38, 6, 57] have been the dominant representation to store 3D information for single scenes and scene/object distributions. Lately, more explicit representations live through a renaissance with 3D Gaussian Splatting [31]. Due to the challenging nature of the task, most methods that enable generalization to 360° novel views only perform reconstruction on the level of individual objects. Even with the recent rise of large-scale generative models, methods that perform extrapolation of larger scenes are still a rarity, often due to missing scalability. In the following, we distinguish between regression-based approaches in Sec. 2.1 and generative approaches in Sec. 2.2.
@@ 2.1 Regression-based Generalizable NVS
Several regression-based models for 3D reconstruction from a few views have been proposed in recent years. One line of work performs generalization over object categories [59, 26, 33, 19, 70] only. On scene level, early methods [76, 13, 41] focus on small-scale setups [27] because of limited capacity or efficiency. Larger scenes require scalable approaches. Image-conditioned neural radiance fields [24] fail in scaling to high resolutions. Image-based rendering methods [1, 45, 49, 71] produce high-quality results in view interpolation but cannot generalize to unseen areas. A related approach is to predict multi-plane images [60, 67, 79, 80], which is limited to small view-point variations only. Multi-view stereo is also a popular way to provide geometry priors for novel view synthesis with deep learning [11, 49, 28]. Moreover, several alternative representations have been introduced, such as neural rays [35], light fields [58, 62], and patches [61].
In contrast to all of the above, we provide high-quality 360° reconstructions of object-centric scenes as well as view inter- and extrapolation on large scenes, given only two input views. Closest to our work are the very recent Splatter Image [63] and pixelSplat [10] following the success of 3D Gaussian Splatting [31] in many domains [14, 72, 18, 23]. In contrast to their purely regression-based approaches, we (1) introduce a semantic feature representation instead of purely explicit Gaussians and (2) model uncertainty explicitly enabling correct generalization to out-of-context views. Thus, we are able to reconstruct full objects in high quality, even if they are only partially observed with two views.
@@ 2.2 Generative Models for NVS
Generative approaches succeed in situations with high uncertainty, i.e. when the conditioning is not sufficient to determine the full reconstruction, or if there is no conditioning at all. A large line of work performs 3D reconstruction of objects with 3D-aware GAN architectures via conditional sampling or inversion [55, 8, 65, 7, 77, 43]. While these methods are able to produce high-quality results, they are not applicable to scenes. Autoregressive transformers have been shown to be able to synthesize novel views that are consistent to some extent with the past sequence of views, but they fail to successfully leverage explicit 3D biases [51, 48]. Following the success of diffusion models [22, 15] for large-scale text-to-image generation [40, 52, 46, 50, 44], there are many approaches to adapting the same concepts for 3D. One line of research trains diffusion models directly on 3D representations such as voxel grids [39, 30, 29], triplanes [56, 12], or point clouds [54, 36, 74], and can implement 3D reconstruction by conditioning or guiding the diffusion process with gradients from reconstructing input views. Similar to 3D-aware GANs, another line [3, 64, 66, 2] integrates rendering of a 3D representation into the denoising architecture for image diffusion to achieve 3D generation and reconstruction purely trained on images. However, both of these approaches do not scale to large scenes due to expensive 3D architectures or slow rendering within the sampling process, respectively. A third group of works [69, 81, 34, 9, 53, 73] extends 2D diffusion models with pose-conditioning for NVS with the additional benefit of compatibility with pre-trained text-to-image generators [50] as strong priors. Closing the loop with regression-based approaches, a particularly effective pose-conditioning proposed by GeNVS [9] is rendering of pixelNeRF [76] features. While scalable to large scenes, these approaches inherit the slow sampling of diffusion models for NVS.
In contrast, our approach is orders of magnitude faster and scales easily to high resolutions due to the efficient Gaussian representation instead of volume rendering, and a lightweight decoder instead of expensive diffusion sampling.
Figure 2: latentSplat architecture. The architecture follows an autoencoder structure. (Left) Two input views are encoded into a 3D variational Gaussian representation using an epipolar transformer and a Gaussian sampling head. (Center) Variational Gaussians allow sampling of spherical harmonics feature coefficients that determine a specific instance of semantic Gaussians. (Right) The sampled instance can be rendered efficiently via Gaussian splatting and a light-weight VAE-GAN decoder.
@@ 3.3 Encoding Reference Views
The variational 3D Gaussian representation described in the previous section is obtained from two given reference views $\mathbf{I}_{1}$, $\mathbf{I}_{2}$ sampled from a video sequence. To this end, we adapt the encoder from pixelSplat [10] to our setting of variational Gaussians by adding the capability of predicting means $\mathbf{h}_{\sigma}$ and variances $\mathbf{h}_{\sigma}$ of spherical harmonic coefficients for view-dependent features of each predicted 3D Gaussian (c.f. Fig 2, left). The encoder consists of three parts: (1) a vision transformer [5], (2) an epipolar transformer [20], and (3) a per-pixel sampling of 3D Gaussians from predicted depth distributions [10], which are shortly outlined in the following. For more details about the encoder architecture, we refer to the supplementals, Charatan et al. [10], and He et al. [20].
@@
With the described encoder, uncertainty is modeled in two ways: First, the Gaussian locations are sampled from the predicted distributions over the rays, modeling uncertainty in 3D Gaussian location. Second, uncertainty in local appearance is modeled via distribution parameter prediction of the variational Gaussians.
@@ 3.5 Training
Reconstruction losses. Similar to LDM [50], we use a combination of a standard L1 loss and LPIPS as a perceptual loss between the decoded feature image $\hat{\mathbf{T}}$ and the target image $\mathbf{T}$:
@@
Auxiliary losses. We further apply an auxiliary loss directly between the color renderings $\hat{\mathbf{T}}_{\mathrm{aux}}$ and the target image $\mathbf{T}$ to provide better gradients to the structural parameters (c.f. Sec 3.2):
@@
Generative loss. To enable correct sampling and generation in uncertain regions, we further optimize our method using GAN losses, adapted from the LDM VAE-GAN decoder [50]. We add and further train a pre-trained discriminator network [25] $D$ from LDM[50] predicting the likelihood of image patches being real. Therefore, our architecture is trained to directly maximize its output:
@@ 4 Experiments
In this section, we detail our experiments made with latentSplat, provide comparisons with state-of-the-art baselines, and verify our architecture design in form of an ablation. Our experiments aim to support the following statements: (1) latentSplat improves on previous methods for two-view interpolation in terms of visual quality, (2) generalizes better to extrapolation, i.e., target views outside of reference views, (3) avoids unnecessary hallucination but sticks to the identity of the observed scene, (4) enables 3D reconstruction by predicting consistent novel views, and (5) maintains the real-time rendering and memory efficiency of 3D Gaussian splatting. We provide additional results in the appendix.
@@ 4.1 Experimental Setup
Datasets. We conduct two-view reconstruction experiments for an object-centric setting and for general video datasets capturing diverse scenes. For the former, we use the Common Objects in 3D [47] dataset, which consists of video captures of real-world objects grouped into categories. Following related work [9], we choose cleaned subsets of hydrants and teddybears, randomly split into 95% training and 5% test data. In all experiments, including baselines, we gradually increase the gap between reference views from initially 8-18 up to 25 frames roughly corresponding to 90° in the 102 frame videos. At the same time, we increasingly randomize the target view selection from pure interpolation until a uniform distribution over all views. For evaluation on general scenes, we leverage RealEstate10k [80], a dataset of home walkthrough clips gathered from about 10000 YouTube videos. We use provided splits and adopt the training curriculum from pixelSplat [10] except of increasing the sampling interval of target views up to 45 frames before and after both reference views to learn extrapolation.
Baselines. We compare our approach against four baselines. pixelNeRF [76] conditions a single NeRF MLP by interpolating pixel-aligned features of reference views. Du et al. [17] is a light field rendering approach that uses multi-view self-attention and cross-attention of target rays to samples along its epipolar lines in the input images. pixelSplat[10] predicts 3D Gaussians along the rays of two reference views that can be efficiently rendered via rasterization. Unlike the previous regression-based approaches, GeNVS [9] proposes a generative diffusion model with view-conditioning via pixelNeRF in a feature space.
@@
Teddybears Ours 53.47 0.004 0.338 0.173 20.83 0.663 71.12 0.010 0.434 0.219 17.71 0.533
Figure 3: Qualitative results on the CO3D dataset [47]. We evaluate two-view NVS on hydrants and teddybears. latentSplat synthesizes high-quality 360° novel views, whereas regression-based approaches suffer from uncertainty resulting in blur.
Metrics. We employ three groups of each two metrics: (1) FID [21] and KID [4] measure the similarity between the distributions of predicted novel views and the corresponding ground truth. They are the established metrics for image synthesis of generative models and reflect visual quality. (2) Perceptual metrics like LPIPS [78] and DISTS [16] leverage features of deep networks for the comparison of images w.r.t. structure and texture. (3) Classical reconstruction metrics such as PSNR and SSIM are still well-established for evaluation of dense-view 3D reconstruction. However, because incorporating plain pixel-wise similarities, these metrics prefer blur over realistic details and are therefore not well-suited for the evaluation of generative methods.
Further details can be found in the supplementary materials and our code.
@@ 4.3 Scene-Level 3D Reconstruction
We report quantitative results for scene-level 3D reconstruction on RealEstate10k in Table 2. Again, latentSplat outperforms all baselines in FID and KID as well as LPIPS and DISTS. Although extrapolation on scene level is a quite different task compared to learning a category-level prior for object-centric videos, we observe the same result that latentSplat generalizes better to extrapolation compared to all baselines, even achieving state-of-the-art in PSNR as well. This highlights the applicability of our method for various kinds of real-world videos. Looking at qualitative examples in Fig. 4, we can see that our approach produces clean and visually pleasing novel views with significantly less artifacts.
@@
Figure 4: Qualitative results on RealEstate10k [80]. We show that in many challenging cases latentSplat outperforms previous methods. This specifically holds for (1) reference views far apart from each other or (2) extrapolation outside of reference views.
4.4 Uncertainty Visualization
We aim to illustrate the uncertainty of our variational Gaussians directly by rendering the standard deviation $\mathbf{h}_{\sigma}$ in Eq. 2, averaged over all feature channels. To deal with empty regions outside of the reference camera frustums, we set the background standard deviation to one, which is in line with our feature map sampling in Eq. 3. The resulting images (second row of Fig. 5(a)) show generally higher uncertainty (dark) for the background, which is either completely invisible or only partly visible in the reference views, compared to the main object, for which the model learns a category-level prior. For the foreground, the model is less certain about details like edges or the fur of teddybears than about plain uniform surfaces, which explains the advantage of the generative decoder w.r.t. a higher level of detail. We provide more examples in the supplementary materials.
4.5 Mesh Reconstruction from Novel Views
High-quality novel view synthesis does not imply strong 3D reconstruction. Especially, considering generative methods, realistic novel views may not be 3D consistent. Therefore, we examine latentSplat’s abilities for downstream 3D reconstruction. Starting with two input images, we sample semantic Gaussians once and render them from all camera poses of the original video. Given these synthesized views, we can employ any surface reconstruction method [68, 75, 23] to obtain textured meshes. We use 2D Gaussian Splatting [23]. The same procedure is applied for the original video and pixelSplat [10] outputs to obtain ground-truth and baseline meshes, respectively. We provide quantitative results for Chamfer distance in Table 3. Our outputs are better suited for reconstruction. This indicates that probabilistic modeling is helpful for faithful surface reconstruction. Qualitative results in Fig. 5(b) further demonstrate that reconstructions with latentSplat are close to the ground truth w.r.t. texture and surface normals.
Table 3: 3D reconstruction on CO3D. Chamfer distance (×$\times 10^{-8}$): Gaussians / mesh from output images versus mesh from ground-truth videos.
Method Hydrants Teddybears
pixelSplat (Gaussians) 49.343 24.965
pixelSplat (mesh from images) 1.815 1.905
Ours (mesh from images) 1.535 1.504
(a) Intermediate results.
(b) Mesh reconstruction.
Figure 5: a) Top to bottom: Given two input images, latentSplat predicts variational Gaussians that model local uncertainty in 3D. Rendering uncertainty reveals a focus on foreground details and out-of-view background. Sampled semantic Gaussians are rendered to features and decoded to final predictions. b) We apply mesh reconstruction on dense synthesized novel views. latentSplat is able to closely approximate the texture and geometry that we can obtain from reconstruction with ground-truth 360° videos.